# Einstein from Noise: Statistical Analysis

### v2

@@ 4 Main Results
We begin by analyzing the regime where $M\to\infty$ and the dimension of the signal $d$ is fixed. In this setting, we show that the Fourier phases of the EfN estimator converge almost surely to those of the underlying template signal, and we characterize the convergence rate. We also analyze the behavior of the Fourier magnitudes. Then, we turn to the high-dimensional regime, where $d\to\infty$. Under additional assumptions, we derive refined asymptotic guarantees for both the phases and magnitudes. Throughout, we assume that the template signal $x\in\mathbb{R}^{d}$ has a unit norm and that its spectrum is non-vanishing, as discussed in the previous section.
Finite-dimensional signal.
We begin with the case where the template signal has a fixed dimension $d$, as captured in the following result, whose proof is provided in Appendix C.
@@
Fix $d\geq 2$ and assume that $\mathsf{X}[k]\neq 0$, for all $10<k\leq d-$ 1.
@@
Theorem 4.1 captures two central properties. The first addresses the convergence of the EfN estimator’s phases to those of the template signal. In addition, the corresponding convergence rate in MSE is proportional to $1/M$. The second result captures the convergence of the EfN estimator’s magnitudes to the term given in the right-hand-side (r.h.s.) of (4.3), which is strictly greater than zero. Thus, the EfN estimator converges to a non-vanishing signal. Interestingly, this term is not necessarily proportional to the magnitudes $\left|\mathsf{X}[k]\right|$ of the template signal.
@@
as $M\to\infty$, where $\mathsf{Q}_{k}$ is a zero-mean Gaussian random variable with variance $\sigma_{\mathsf{Q}}^{2}[k]=C_{k}/M$, and the constant $C_{k}$ admits a closed-form expression. By leveraging properties of cyclo-stationary Gaussian processes, which is justified by the circulant structure of the problem, we establish that $C_{k}<\infty$ for all $10\leq k\leq d-$ 1. This directly leads to the results stated in (4.1)–(4.2).
High-dimensional regime.
We now turn to the high-dimensional setting where $d\to\infty$, taken after the limit $M\to\infty$. In this regime, we impose additional technical conditions on the template signal, formalized in Assumption 4.2. Intuitively, these conditions reflect the empirical phenomenon illustrated in Figure 3, where a flatter PSD, which corresponds to a more rapidly decaying autocorrelation function, results in a higher alignment between the template and the estimator.
More precisely, Assumption 4.2 requires control over the decay of both the autocorrelation function and the spectral magnitudes as functions of $d$. Specifically, the autocorrelation $\mathsf{R}_{\mathsf{X}\mathsf{X}}$, defined as the Fourier transform of PSD $|\mathsf{X}|^{2}$, must decay faster than $1/\log d$, and the maximum magnitude among nonzero Fourier components $|\mathsf{X}[k]|$ must decay faster than $1/\sqrt{\log d}$. In addition, we assume the DC component is vanishing, i.e., $|\mathsf{X}[0]|=0$, to avoid degeneracies in alignment.
Assumption 4.2.
Let $\mathsf{X}={\cal F}\left\{x\right\}$ and let $\mathsf{R}_{{\mathsf{X}\mathsf{X}}}$ be the auto-correlation of the signal $x$. We say that the template signal $x$ satisfies Assumption 4.2 if the following hold:
@@
    The auto-correlation satisfies,
    $\lim_{d\to\infty}\mathsf{R}_{{\mathsf{X}\mathsf{X}}}[d]\cdot\log d=0.$ (4.4)
@@
    The magnitudes satisfy,
    $\lim_{d\to\infty}\left\{{\underset{0<k\leq d-1}{\max}\left\{\left|\mathsf{X}[k]\right|\right\}}\cdot\sqrt{\log d}\right\}=0.$ (4.5)
@@
    The signal’s DC component is zero, i.e., ${\left|\mathsf{X}[0]\right|}=0$.
Although the conditions in Assumption 4.2 may seem technical, they are essential for establishing Theorem 4.3, which relies on classical limit theorems for the maxima of stationary Gaussian processes, most notably, convergence to the Gumbel distribution [26, 10, 2, 4]. Each part of the assumption plays a specific role: Part (1) ensures that the noise process lacks long-range dependencies, which corresponds to a sufficiently flat PSD; Part (2) guarantees that no individual Fourier component dominates the behavior of the EfN estimator. The final condition, requiring the DC component to vanish (i.e., $|\mathsf{X}[0]|=0$), is not strictly necessary from an empirical standpoint but is introduced to streamline the theoretical analysis.
@@
Assume that $\mathsf{X}[k]\neq 0$, for all $10<k\leq d-$ 1, and that $x$ satisfies Assumption 4.2. Then,
@@
    For any $10\leq k\leq d-$ 1, we have,
    $\lim_{d\to\infty}\lim_{M\to\infty}\frac{\mathbb{E}|\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]|^{2}}{1/(M\log d)}\frac{1}{1/(4\left|\mathsf{X}[k]\right|^{2})}=1.$ (4.6)
@@
    For any $10\leq k\leq d-$ 1, we have,
    $\frac{1}{2\sigma\sqrt{\log d}}\frac{|\mathsf{\hat{X}}[k]|}{\left|\mathsf{X}[k]\right|}\xrightarrow[]{\mathsf{a.s.}}1,$ (4.7)
@@
The proof of Theorem 4.3 is presented in Appendix F. Based on Theorem 4.3, as $M,d\to\infty$, the convergence rate of the Fourier phases of the EfN estimator is inversely proportional to the Fourier magnitude square. In addition, unlike Theorem 4.1, the Fourier magnitudes of the EfN estimator converge to those of the template signal, up to a constant factor. Therefore, when $d\to\infty$, under Assumption 4.2, the signal $\hat{x}$ recovers the template signal up to a known normalization factor. This recovery, in turn, implies that the normalized cross-correlation between the template and the EfN estimator approaches unity. To obtain the refined convergence rate in Theorem 4.3 for $d\to\infty$ (namely, that $C_{k}=\frac{1}{4\left|\mathsf{X}[k]\right|^{2}\log d}$), we utilize results from the theory of extrema of Gaussian processes, particularly, the convergence of the maximum of a stationary Gaussian process to the Gumbel distribution, see, e.g., [26, 10, 2, 4].
@@
Figure 4: Comparison between analytic expression and Monte-Carlo simulations for high-dimensional signals, $d$, and for signals with varying power spectral densities. The analytic expressions for the Fourier phases convergence and Fourier magnitudes are given by (4.6) and (4.7), respectively. (a) Three signals with varying PSDs are examined. (b) Monte-Carlo simulation of the Pearson cross-correlation between the template, $x$, and the estimator, $\hat{x}$. As $d$ increases, the correlation between the estimator and the template vector increases, especially for signals with faster-decaying auto-correlations. (c) The analytic expression (4.6) closely aligns with Monte-Carlo simulations and shows better agreement for signals characterized by faster-decaying auto-correlations. The simulations were conducted with $d=8192$ and $M=10000$, based on 2000 Monte-Carlo trials.
@@ 5.1 Positive Correlation
Let $d\geq 2$, and suppose the observations $\{y_{i}\}_{i=0}^{M-1}$ are drawn i.i.d. according to the model in (5.1). Let $x\in\mathbb{R}^{d}$ denote the template signal, and assume its discrete Fourier transform $\mathsf{X}$ satisfies $\mathsf{X}[k]\neq 0$ for all $11\leq k\leq d-$ 1. Let $\hat{x}$ be the EfN estimator computed from the observations $\{y_{i}\}$. Then, as $M\to\infty$, the following inequality holds almost surely,
@@
The proof of Proposition 5.1 is provided in Appendix G. This result implies that the EfN estimator is positively correlated with the true template signal. Although this is a weaker guarantee than the Fourier phase convergence obtained under Gaussian white noise, it still ensures that the estimator retains meaningful structural information from the template.
@@
The proof of Theorem 5.2 is given in Appendix H. In essence, this result extends the Fourier phase convergence of Theorem 4.1 to a broader class of noise distributions in the high-dimensional setting. The main idea of the proof is to apply the functional central limit theorem to the DFT coefficients [34, 12, 11]. As $d\to\infty$, the Fourier components of the noise converge in distribution to those of a circulant Gaussian random process, owing to the i.i.d. structure of the entries in $y_{i}$. This asymptotic Gaussianity enables us to apply the same analytical framework developed for the white noise case to establish convergence of the Fourier phases.
@@
Figure 5: The impact of noise statistics and signal dimension ($d$) on Fourier phase convergence. Each panel displays the mean squared error (MSE) between the Fourier phases of the true template and those estimated by EfN, shown for three representative Fourier components. The dashed line represents the theoretical $1/M$ convergence rate. Columns correspond to different noise distributions: white Gaussian noise, i.i.d. noise drawn from a uniform distribution over the interval $[0,1]$, and i.i.d. Poisson noise with parameter $\lambda=10$. Rows correspond to increasing signal dimensions: $d=8$, $32$, and $1024$. For white Gaussian noise, the Fourier phases converge at the expected $1/M$ rate across all signal dimensions, in agreement with Theorem 4.1. In contrast, under uniform and Poisson noise, the MSE plateaus at low dimensions. However, increasing the signal dimension restores convergence, even under non-Gaussian noise, consistent with the high-dimensional regime described in Theorem 5.2. Notably, for $d=1024$, all three noise models produce similar MSE values across the selected Fourier components, suggesting that their phase noise statistics become nearly indistinguishable. Each data point represents an average of 300 Monte Carlo trials.
@@ 5.3 Circulant Gaussian process
Let $d\geq 2$ be fixed, and suppose the observations $\{y_{i}\}_{i=0}^{M-1}$ are i.i.d samples drawn from the multivariate normal distribution $\mathcal{N}(0,\Sigma)$, where $\Sigma$ is a symmetric circulant matrix as defined in Definition 5.3. Assume further that the eigenvalues of $\Sigma$ are strictly positive, and that the template signal $x\in\mathbb{R}^{d}$ satisfies $\mathsf{X}[k]\neq 0$ for all $11\leq k\leq d-$ 1. Let $\hat{x}$ denote the EfN estimator under this noise model. Then, for each $10\leq k\leq d-$ 1:
@@
The proof of Proposition 5.4 is given in Appendix I. In essence, this result serves as a generalization of Theorem 4.1, which considered the case of white Gaussian noise, to the broader setting of symmetric circulant Gaussian noise. Notably, white noise with covariance $\sigma^{2}I_{d\times d}$ is a special case of circulant noise, making this extension a natural generalization. The critical insight here is that circulant covariance matrices remain diagonalizable in the Fourier basis, which preserves the independence of the DFT coefficients and enables phase convergence to proceed as in the white Gaussian case.
@@
Figure 6: The impact of the noise covariance structure and signal dimension ($d$) on Fourier phase convergence. Each plot shows the mean squared error (MSE) between the Fourier phases of the ground-truth template and those estimated by EfN, evaluated across three spectral components. The dashed line indicates the theoretical $1/M$ convergence rate. Columns correspond to three types of noise: (1) white Gaussian noise with covariance $\Sigma=\sigma^{2}I_{d\times d}$, (2) symmetric circulant covariance, and (3) a Toeplitz covariance matrix that is not circulant. Rows represent increasing signal dimensions: $d=8$, $32$, and $1024$. Under white Gaussian noise, the Fourier phases converge at the expected $1/M$ rate, independent of the signal dimension (Theorem 4.1). A similar trend is observed when the noise has a circulant covariance structure: the same $1/M$ scaling holds, although the MSE is different compared to the white noise case. In contrast, for a Toeplitz covariance matrix that is not circulant, the MSE plateaus at small signal dimensions, indicating a failure of convergence. However, when the signal dimension increases to $d=1024$, convergence at the $1/M$ rate is restored even under this more structured noise model. Each data point represents an average of 300 Monte Carlo trials.

### v3

@@ 4 Main Results
We begin by analyzing the regime where $M\to\infty$ and the dimension of the signal $d$ is fixed. In this setting, we show that the Fourier phases of the EfN estimator converge almost surely to those of the underlying template signal, and we characterize the convergence rate. We also analyze the behavior of the Fourier magnitudes. Then, we turn to the high-dimensional regime, where $d\to\infty$. Under additional assumptions, we derive refined asymptotic guarantees for both the phases and magnitudes. Throughout, we assume that the template signal $x\in\mathbb{R}^{d}$ has a unit norm and that its spectrum is non-vanishing, i.e., $\mathsf{X}[k]\neq 0$ for every $0<k\leq d-1$, as discussed in the previous section.
4.1 Finite-dimensional signal
We begin with the case where the template signal has a fixed dimension $d$, as captured in the following result, whose proof is provided in Appendix B.5.
@@ 4.1 Finite-dimensional signal
Fix $d\geq 2$ and assume that $\mathsf{X}[k]\neq 0$, for all $0<k\leq d-1$.
@@
Theorem 4.1 captures two central properties. The first addresses the convergence of the EfN estimator’s phases to those of the template signal. In addition, the corresponding convergence rate in MSE is proportional to $1/M$. The second result captures the convergence of the EfN estimator’s magnitudes to the term given in the right-hand-side (r.h.s.) of (4.3), which is strictly greater than zero. Thus, the EfN estimator converges to a non-vanishing signal. Interestingly, this term is not necessarily proportional to the magnitudes $\left|\mathsf{X}[k]\right|$ of the template signal and, thus, the EfN estimator reproduces (asymptotically) only the phases of the template signal but not the magnitudes.
@@
as $M\to\infty$, where $\mathsf{Q}_{k}$ is a zero-mean Gaussian random variable with variance $\sigma_{\mathsf{Q}}^{2}[k]=C_{k}/M$, and the constant $C_{k}$ admits a closed-form expression. By leveraging properties of cyclo-stationary Gaussian processes, which is justified by the circulant structure of the problem, we establish that $C_{k}<\infty$ for all $0\leq k\leq d-1$. This directly leads to the results stated in (4.1)–(4.2).
4.2 High-dimensional regime
We now turn to the high-dimensional setting where $d\to\infty$, taken after the limit $M\to\infty$. In this regime, we impose additional technical conditions on the template signal, formalized in Assumption 4.2. Intuitively, these conditions reflect the empirical phenomenon illustrated in Figure 3, where a flatter PSD, which corresponds to a more rapidly decaying autocorrelation function, results in an improved alignment between the template and the estimator.
More precisely, Assumption 4.2 below requires control over the decay of both the autocorrelation function and the spectral magnitudes as functions of $d$. Specifically, for a length-$d$ signal $x\in\mathbb{R}^{d}$, we define the (circular) autocorrelation in the time domain by
$\mathsf{R}_{\mathsf{X}\mathsf{X}}[\tau]\triangleq\frac{1}{\sqrt{d}}\sum_{n=0}^{d-1}{x}[n]\;{x}[n+\tau\!\!\!\pmod{d}$], (4.4)
for $\tau\in\{0,1,\ldots,d-1\}$. By the discrete Wiener-Khinchin theorem, this is equivalent to taking the inverse discrete Fourier transform of the PSD, namely,
$\mathsf{R}_{\mathsf{X}\mathsf{X}}[\tau]=\mathcal{F}^{-1}\!\left\{\,|\mathsf{X}|^{2}\,\right\}[\tau]=\frac{1}{\sqrt{d}}\sum_{k=0}^{d-1}|\mathsf{X}[k]|^{2}\,e^{j\frac{2\pi}{d}k\tau},$ (4.5)
where $\mathsf{X}=\mathcal{F}\{\mathsf{x}\}$ is the DFT of $x$. Assumption 4.2 requires that the autocorrelation $\mathsf{R}_{\mathsf{X}\mathsf{X}}[\tau]$ decay faster than $1/\log d$, and that the maximum magnitude among nonzero Fourier components $\max_{k\neq 0}|\mathsf{X}[k]|$ decay faster than $1/\sqrt{\log d}$. In addition, we assume the DC component is vanishing, i.e., $|\mathsf{X}[0]|=0$, to avoid degeneracies in alignment.
Assumption 4.2 (High-dimensional regularity of the template).
Consider a sequence of template signals $\{x^{(d)}\}_{d\in\mathbb{N}}$ with $x^{(d)}\in\mathbb{R}^{d}$, and let $\mathsf{X}^{(d)}=\mathcal{F}\{x^{(d)}\}$ denote their DFT. Let $\mathsf{R}^{(d)}_{\mathsf{X}\mathsf{X}}$ denote the autocorrelation of $x^{(d)}$, as defined in (4.4). When taking the limit $d\to\infty$, we assume the signals are normalized, namely $\|x^{(d)}\|_{2}=1$ for all $d\in\mathbb{N}$. We say that the template sequence $\{x^{(d)}\}_{d\in\mathbb{N}}$ satisfies Assumption 4.2 if the following hold:
@@ 4.2 High-dimensional regime
    Autocorrelation decay. The autocorrelation away from the zero lag satisfies
    $\lim_{d\to\infty}\left(\max_{1\leq\tau\leq d-1}\big|\mathsf{R}^{(d)}_{\mathsf{X}\mathsf{X}}[\tau]\big|\right)\cdot\log d=0.$ (4.6)
@@
    Spectral magnitude decay. The non-DC Fourier magnitudes satisfy
    $\lim_{d\to\infty}\left(\max_{1\leq k\leq d-1}\big|\mathsf{X}^{(d)}[k]\big|\right)\cdot\sqrt{\log d}=0.$ (4.7)
@@
    Vanishing DC component. The signal’s DC component is zero, i.e., ${\left|\mathsf{X}^{(d)}[0]\right|}=0$.
Although the conditions in Assumption 4.2 may seem technical, they are essential for establishing Theorem 4.3, which relies on classical limit theorems for the maxima of stationary Gaussian processes, most notably, convergence to the Gumbel distribution [29, 12, 2, 4]. Each part of the assumption plays a specific role: Part (1) ensures that the noise process lacks long-range dependencies, which corresponds to a sufficiently flat PSD; Part (2) guarantees that no individual Fourier component dominates the behavior of the EfN estimator. The final condition, requiring the DC component to vanish (i.e., $|\mathsf{X}^{(d)}[0]|=0$), is not strictly necessary from an empirical standpoint but is introduced to streamline the theoretical analysis.
@@
Assume that $\mathsf{X}^{(d)}[k]\neq 0$, for all $0<k\leq d-1$, and that $x$ satisfies Assumption 4.2. Then,
@@
    For any $0\leq k\leq d-1$, we have,
    $\lim_{d\to\infty}\lim_{M\to\infty}\frac{\mathbb{E}|\phi_{\mathsf{\hat{X}}^{(d)}}[k]-\phi_{\mathsf{X}^{(d)}}[k]|^{2}}{1/(M\log d)}\frac{1}{1/(4\left|\mathsf{X}^{(d)}[k]\right|^{2})}=1.$ (4.8)
@@
    For any $0\leq k\leq d-1$, we have,
    $\frac{1}{\sigma\sqrt{2\log d}}\frac{|\mathsf{\hat{X}}^{(d)}[k]|}{\left|\mathsf{X}^{(d)}[k]\right|}\xrightarrow[]{\mathsf{a.s.}}1,$ (4.9)
@@
The proof of Theorem 4.3 is presented in Appendix D.3. Based on Theorem 4.3, as $M,d\to\infty$, the convergence rate of the Fourier phases of the EfN estimator is inversely proportional to the squared Fourier magnitude. In addition, unlike the fixed-$d$ result in Theorem 4.1, the high-dimensional regime reveals an explicit dependence on $d$ in the phase-error constant. Specifically, while Theorem 4.1 states that for each fixed dimension $d$, (4.2) holds with a finite constant $C_{k}$ (which depends on $d$ and on the template), Theorem 4.3 makes this dependence explicit when $d\to\infty$: comparing (4.8) with (4.2) shows that in the high-dimensional limit the constant scales as
$C_{k}=\frac{1}{4\,|\mathsf{X}^{(d)}[k]|^{2}\,\log d}.$ (4.10)
Intuitively, the $\log d$ factor arises from the alignment step: $\hat{\mathsf{R}}_{i}$ is chosen by maximizing a stationary Gaussian correlation process over $d$ shifts, and the maximizer is governed by extreme-value statistics. In particular, the maximum of such a process grows on the scale $\sqrt{2\log d}$; see, e.g., [29, 12, 2, 4], and the proof in Appendix D.
Moreover, unlike Theorem 4.1 for a fixed $d$, Theorem 4.3 also shows that the Fourier magnitudes of the EfN estimator satisfy (4.9), namely they recover the template magnitudes up to the known normalization factor $\sigma\sqrt{2\log d}$. Therefore, when $d\to\infty$ under Assumption 4.2, the normalized estimate $\hat{x}^{(d)}$ recovers the template signal, which in turn implies that the normalized cross-correlation between the template and the EfN estimator approaches unity.
@@
Figure 4: Comparison between analytic expression and Monte-Carlo simulations for high-dimensional signals, $d$, and for signals with varying power spectral densities. The analytic predictions for Fourier-phase convergence and Fourier-magnitude scaling are given by (4.8) and (4.9), respectively. (a) Template PSDs for three template families at a representative dimension $d=8192$. For each dimension $d$, the template $x^{(d)}\in\mathbb{R}^{d}$ is generated directly at length $d$ as an exponentially decaying signal, $x^{(d)}_{\ell}[m]\propto\exp(-m/\alpha_{\ell})$, $m=0,1,\ldots,d-1$, with decay parameters $\alpha_{\ell}\in\{0.02,\,2,\,10\}$ (Signals 1-3, respectively), followed by mean removal and normalization. (b) Monte-Carlo estimates of the Pearson cross-correlation $\mathrm{PCC}(x^{(d)}_{\ell},\hat{x}^{(d)}_{\ell})$ between the template $x^{(d)}_{\ell}$ and the EfN estimate $\hat{x}^{(d)}_{\ell}$ as a function of the signal length $d$ (with fixed sample size $M=10^{4}$). As $d$ increases, the correlation increases, particularly for templates with slower-decaying PSDs. (c) Per-frequency phase mean-squared error $\mathbb{E}|\phi_{\hat{\mathsf{X}}^{(d)}_{\ell}}[k]-\phi_{\mathsf{X}^{(d)}_{\ell}}[k]|^{2}$ at $d=8192$: Monte-Carlo estimates (blue) are compared with the asymptotic expression (red), i.e., the large-$(M,d)$ closed-form approximation predicted by (4.8). All Monte-Carlo curves are averaged over 20002000 independent trials.
@@ 5.1 Positive correlation
Let $d\geq 2$, and suppose the observations $\{y_{i}\}_{i=0}^{M-1}$ are drawn i.i.d. according to the model in (5.1). Let $x\in\mathbb{R}^{d}$ denote the template signal, and assume its discrete Fourier transform $\mathsf{X}$ satisfies $\mathsf{X}[k]\neq 0$ for all $1\leq k\leq d-1$. Let $\hat{x}$ be the EfN estimator computed from the observations $\{y_{i}\}$. Then, as $M\to\infty$, the following inequality holds almost surely,
@@
The proof of Proposition 5.1 is provided in Appendix E. This result implies that the EfN estimator is positively correlated with the true template signal. Although this is a weaker guarantee than the Fourier phase convergence obtained under Gaussian white noise, it still ensures that the estimator retains meaningful structural information from the template.
@@
The proof of Theorem 5.2 is given in Appendix F. In essence, this result extends the Fourier phase convergence of Theorem 4.1 to a broader class of noise distributions in the high-dimensional setting. The main idea of the proof is to apply the functional central limit theorem to the DFT coefficients [37, 14, 13]. As $d\to\infty$, the Fourier components of the noise converge in distribution to those of a circulant Gaussian random process, owing to the i.i.d. structure of the entries in $y_{i}$. This asymptotic Gaussianity enables us to apply the same analytical framework developed for the white noise case to establish convergence of the Fourier phases.
@@
Figure 5: The impact of noise statistics and signal dimension ($d$) on Fourier phase convergence. Each panel displays the mean squared error (MSE) between the Fourier phases of the true template and those estimated by EfN, shown for three representative Fourier components. The dashed line represents the theoretical $1/M$ convergence rate. Columns correspond to different noise distributions: white Gaussian noise, i.i.d. noise drawn from a uniform distribution over the interval $[0,1]$, and i.i.d. Poisson noise with parameter $\lambda=10$. Rows correspond to increasing signal dimensions: $d=8$, 3232, and $1024$. For white Gaussian noise, the Fourier phases converge at the expected $1/M$ rate across all signal dimensions, in agreement with Theorem 4.1. In contrast, under uniform and Poisson noise, the MSE plateaus at low dimensions. However, increasing the signal dimension restores convergence, even under non-Gaussian noise, consistent with the high-dimensional regime described in Theorem 5.2. Notably, for $d=1024$, all three noise models produce similar MSE values across the selected Fourier components, suggesting that their phase noise statistics become nearly indistinguishable. Each data point represents an average of 300 Monte Carlo trials.
@@ 5.3 Circulant Gaussian process
Let $d\geq 2$ be fixed, and suppose the observations $\{y_{i}\}_{i=0}^{M-1}$ are i.i.d samples drawn from the multivariate normal distribution $\mathcal{N}(0,\Sigma)$, where $\Sigma$ is a symmetric circulant matrix as defined in Definition 5.3. Assume further that the eigenvalues of $\Sigma$ are strictly positive, and that the template signal $x\in\mathbb{R}^{d}$ satisfies $\mathsf{X}[k]\neq 0$ for all $1\leq k\leq d-1$. Let $\hat{x}$ denote the EfN estimator under this noise model. Then, for each $0\leq k\leq d-1$:
@@
The proof of Proposition 5.4 is given in Appendix G. In essence, this result serves as a generalization of Theorem 4.1, which considered the case of white Gaussian noise, to the broader setting of symmetric circulant Gaussian noise. Notably, white noise with covariance $\sigma^{2}I_{d\times d}$ is a special case of circulant noise, making this extension a natural generalization. The critical insight here is that circulant covariance matrices remain diagonalizable in the Fourier basis, which preserves the independence of the DFT coefficients and enables phase convergence to proceed as in the white Gaussian case.
@@
Figure 6: The impact of the noise covariance structure and signal dimension ($d$) on Fourier phase convergence. Each plot shows the mean squared error (MSE) between the Fourier phases of the ground-truth template and those estimated by EfN, evaluated across three spectral components. The dashed line indicates the theoretical $1/M$ convergence rate. Columns correspond to three types of noise: (1) white Gaussian noise with covariance $\Sigma=\sigma^{2}I_{d\times d}$, (2) symmetric circulant covariance, and (3) a Toeplitz covariance matrix that is not circulant. Rows represent increasing signal dimensions: $d=8$, 3232, and $1024$. Under white Gaussian noise, the Fourier phases converge at the expected $1/M$ rate, independent of the signal dimension (Theorem 4.1). A similar trend is observed when the noise has a circulant covariance structure: the same $1/M$ scaling holds, although the MSE is different compared to the white noise case. In contrast, for a Toeplitz covariance matrix that is not circulant, the MSE plateaus at small signal dimensions, indicating a failure of convergence. However, when the signal dimension increases to $d=1024$, convergence at the $1/M$ rate is restored even under this more structured noise model. Each data point represents an average of 300 Monte Carlo trials.